OpenStudy (anonymous):

WILL REWARD! Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

@johnweldon1993 please help"?

OpenStudy (johnweldon1993):

First, what do we know about slopes of parallel lines?

OpenStudy (anonymous):

well parrallel lines never touch and go the same way

OpenStudy (johnweldon1993):

Right, in other words, they have the same slope If the slope was ANY different, they would eventually touch correct?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

So in your equation up there \[\large y = -3x + 12\] What is the slope?

OpenStudy (anonymous):

so the slope would be -3x?

OpenStudy (johnweldon1993):

Well the slope would be just -3 \(\large y = mx + b\) where 'm' is the slope

OpenStudy (johnweldon1993):

So there, we know the slope of the line that we need. And we also have a point along that line (-1 , 6) Since we have a point and a slope, we can use the point slope formula \[\large y - y_0 = m(x - x_0)\]

OpenStudy (anonymous):

so y-6=-3(x+1)

OpenStudy (johnweldon1993):

Replace \(\large y_0\) with your y-coordinate And \(\large x_0\) with your x coordinate and finally \(\large m\) with your slope

OpenStudy (anonymous):

am i right?

OpenStudy (johnweldon1993):

Completely correct, \[\large y - 6 = -3(x + 1)\] Now just distribute out the right hand side, and then add 6 to both sides to isolate and solve for y

OpenStudy (anonymous):

y=-3x+3

OpenStudy (johnweldon1993):

There you go!

OpenStudy (anonymous):

but thats not one of the choices...

OpenStudy (johnweldon1993):

What are your choices?

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

-1/3x means -one thirdx

OpenStudy (johnweldon1993):

That is a completely different question...

OpenStudy (anonymous):

omg im soooo sorry! i had the on i typed earlier D:

OpenStudy (anonymous):

can you help still again im sooo sorry!

OpenStudy (johnweldon1993):

lol no problem! But since you know the process now can you figure this one out?

OpenStudy (anonymous):

woould the m=-3? thats one of the things tha confuses me

OpenStudy (anonymous):

ik that y=-2 and x=4

OpenStudy (johnweldon1993):

Nope, positive 3 So we start off with \[\large 3x - y = -3\] Add y to both sides, then add 3 to both sides \[\large y = 3x + 3\] you could have alternatively subtracted 3x from both sides, then divided everything by '-' to solve for y

OpenStudy (anonymous):

thats something i didnt know thanks!.

OpenStudy (anonymous):

so y+2=3(x-4)?

OpenStudy (johnweldon1993):

Yeah you always want to solve for 'y' it is the easiest way (in my own opinion that is) So yes that is correct, and simplify it a bit more...

OpenStudy (anonymous):

k so i distribute the 3 on the right side, add 2 on both sides...

OpenStudy (anonymous):

y=3x-14

OpenStudy (johnweldon1993):

If you meant to say "subtract 2 from both sides" then yes

OpenStudy (anonymous):

thank you soo much john!

OpenStudy (johnweldon1993):

Anytime!

OpenStudy (anonymous):

i turned it in and got an 1 if it weren't for you i wouldve gotten a c you a life saver!

OpenStudy (anonymous):

A*

OpenStudy (johnweldon1993):

Well I'm glad! If you need anymore help feel free to message me

OpenStudy (anonymous):

ok thanks